OpenStudy (anonymous):

Solve the system of linear equations below. y = -2x + 19 y = x + 7

OpenStudy (anonymous):

Substitute the 2nd equation into the first equation x+7=-2x+19

OpenStudy (anonymous):

i got x = -2x + 12

OpenStudy (anonymous):

x+7=-2x+19 3x = 12 x=4 To get y, substitute 4 into the eq above to get y 4+7=y 11=y check to be sure both eq work -2(4) +19 =y -8+19 =11 solution is always an ordered pair (4,11) Knowing that will make your teacher happy

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

good luck